OpenStudy (anonymous):

The table below shows the squares of different numbers: Number 2 -2 3 -3 (x) Square of the number 4 4 9 9 (y) Part A: Does the table represent y as a function of x? Justify your answer. (5 points) Part B: The total cost f(x), in dollars, for renting a rowboat for x hours is shown below: f(x) = 20 + 3x What is the value of f(150), and what does f(150) represent? (5 points)

OpenStudy (anonymous):

@TuringTest @ganeshie8

OpenStudy (acxbox22):

y is a function of x because every input value(x) will have only one outcome(y) f(150)=470 , it means that renting a rowboat for 150 hours will cost $450 +$20 for a flat fee

OpenStudy (anonymous):

Are you positive that B. is correct?

OpenStudy (turingtest):

do not solicit plain answers. use what you were told to make a judgement, or ask more questions to further your understanding, otherwise it appears as though you are trying to get the answer without learning, which is not the point of this site.

OpenStudy (anonymous):

Ok. I actually do understand part A. I was just checking my answer. :) And, of course I will check the answers before submitting them.

OpenStudy (anonymous):

Um, @TuringTest would you mind explaining in like 2 sentences about part A so I can understand it a little more?

OpenStudy (acxbox22):

true but most likely B. is right because f(150)=20+3(150) f(150)=20+450 f(150)=470

OpenStudy (anonymous):

Alright, got it. I did the problem as well, and it makes sense. :) Thank you.

OpenStudy (acxbox22):

no problem

OpenStudy (anonymous):

@acxbox22 um, one more question. What does f(150) represent?

OpenStudy (anonymous):

Is it dollars or hours?

OpenStudy (turingtest):

consider the relation between each pair of (x, y) (2, 4) (-2, 4) (3, 9) (-3, 9) in each case y=f(x)=x^2, that is, put in any x into the function f(x)=x^2 and you will get y. This indicates y is a function of x.

OpenStudy (anonymous):

Ok, I understand. :) One more question, what would f(150) represent? I'm thinking it represents the amount of money the people who rented the rowboat owe after a certain amount of hours.

OpenStudy (acxbox22):

f(150) represents that you rent the boat for 150 hours which is a long time. So if you rent the boat for 150 hours you owe $470 to the people you rented the boat from.

OpenStudy (anonymous):

Thank you. :)